Test the claim that the mean GPA of night students is significantly different than the mean GPA of day students at the 0.1 significance level.

The null and alternative hypothesis would be:

H0:pN≥pDH0:pN≥pD
H1:pN<pDH1:pN<pD

H0:pN≤pDH0:pN≤pD
H1:pN>pDH1:pN>pD

H0:μN≥μDH0:μN≥μD
H1:μN<μDH1:μN<μD

H0:pN=pDH0:pN=pD
H1:pN≠pDH1:pN≠pD

H0:μN≤μDH0:μN≤μD
H1:μN>μDH1:μN>μD

H0:μN=μDH0:μN=μD
H1:μN≠μDH1:μN≠μD



The test is:

left-tailed

two-tailed

right-tailed



The sample consisted of 65 night students, with a sample mean GPA of 3.35 and a standard deviation of 0.04, and 65 day students, with a sample mean GPA of 3.33 and a standard deviation of 0.06.

The test statistic is:  (to 2 decimals)

The p-value is:  (to 2 decimals)
Answer #1

1)

H0:μN=μD

H1:μN≠μD

two-tailed test (since we are not looking difference in any particular direction)

2)

since sample size is large, we can use standard normal distribution.

std error σx1-x2=√(σ21/n122/n2)    = 0.009
test stat z =(x1-x2o)/σx1-x2     = 2.24
p value : = 0.03